hydrogen peroxide is a powerful exciting agent used in concentrated solution in rocket fuels and in diluted solution is hair bleach. An aqueous solution of 30% (m/m) has a density of 1.11 g/m. calculate the concentration in terms of mole fraction of hydrogen peroxide and molarity

1
Expert's answer
2022-02-04T08:56:02-0500

Molality= m(solute)/m(solvent). p=m/V. Molality = 30/0.07= 428,571429.